Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In the midst of the move from D to DnC; closed out D config well (see later). Ongoing work with both 3-bit and 4-band testing this week.
  • Key dates:
    • 13 Sep: Begin move to DnC configuration
    • 13-22 Sep: 74 MHz array testing campaign
    • 15 Sep: Call for Proposals
      • Note: Ku/X-wide will not be offered in main call (allowed for exploratory).
    • 01 Oct: Proposal deadline (A config; lowband for end of B config)
    • 04 Oct: Begin move to C configuration
    • 13-15 Oct: Phased Array testing campaign
  • Staff Travel:
    • Chandler: 01 Sep - 20 Sep
    • Rupen: 11 Sep - 21 Sep
    • Sowinski: 20 Sep - 15 Oct
    • McMullin: 01-10 Oct (vacation/Chile)
    • van Moorsel: 11-15 Oct
    • Wrobel: 24 Oct-27 Oct
    • DRAO: Dougherty: 11 Oct - 12 Oct (SKA; ABQ)
  • Visitors:
    • Brent Carlson: 18 October - 25 October; second trip post SKA meeting TBD
    • DRAO: Dougherty: 12 Oct - 17 Oct (NRAO-DSOC)
    • Adam Deller/Joe Lazio: 13-15 Oct (phased array testing)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Notes: 5 antennas moving (EA08, EA20, EA18, EA13, EA10); working on baselines/delays; Get L (EA06),C (EA14), S (EA14) band receivers this week.
      • Antennas with 3-bit samplers: EA15, EA22, EA27, EA28
      • Antennas with 4-band dipoles: EA01(West), EA03(East), EA06(North), EA11(East), EA19(West), EA28(North)
    • Links:
  • WIDAR (Rupen):
    • Notes: Switch installation on Wednesday of this week. Report on 7 vs 8 node observing; Miriam reviewed the comparison OSRO data sets; no major issues with zeros; other problems with the data in phase drifts, etc that are being explored. More on RSRO + missing/mislabled in minutes.

3-bit testing status report (Perley)

  • Notes: See attached plots as shown during the meeting (incremental views of calibration); but see more noise in RCP; 20% higher (beyond expectation from quantization loss between 3-bit and 8-bit sampling). EA22 sampler is very bad; pulled out of the antenna; just repaired (FPGA transplant) and will go back in shortly. Note we have missed our 'drop-dead' date for the order of the production equipment (which will cause delays in the availability for the widest bandwidths for RSRO); problem continues to be explored at high priority.
  • 2GHZ-15X22-CAL.ps: 2GHZ-15X22-CAL.ps
  • 8B-LCP.ps: 8B-LCP.ps
  • 3B-LCP.ps: 3B-LCP.ps
  • 2GHZ-15X22-RAW.ps: 2GHZ-15X22-RAW.ps

Software status and updates (Butler)

  • Links:
  • Notes: CBE Node 6 issues, under analysis; parameter database went down; not yet understood. 2 hours of downtime this week on Tuesday for the archive retrieval. Push on CBE time averaging this week for software development. Switched power iteration this week for creation and storing in the SDM to filler. SM: plan for testing? RP: switch 1dB in and out to see stability; also long observation over changing conditions (airmasses at K band); will do once the basic mechanics are firmly established. SM has a plotter for viewing the existing SysPower table; Eric/Bill are working on the AIPS/UVFITS implementation.
  • CASA Update (new functionality available/issues/enhancements). Development planning for next release (next week). Added parameter for channelized setjy application (not a single value per spectral window).
  • AIPS Update (new functionality available/issues/enhancements). Noted that y2k performance task exposed a bug in imaging (not including components from some facets); being worked on.

Observing recommendations update (Ops)

Commissioning priorities for the week (McMullin, Chandler)

  • 1) Configuration move; re-establishing antenna operations
  • 2) Software testing and updates (CBE, CM; work toward CBE averaging, work on missing/mis-labeled, robustness)
  • 3) Science observing as possible (Any array projects)
  • Interrupt status for additional 3-bit testing, 4-band tests needed.

Commissioning Reports

Spectral Line: Momjian: Zeeman calibration test status.

    • See: https://bugs.aoc.nrao.edu/browse/ECSV-105; in progress. Looking at 6.7GHz CH3OH maser (see figures in JIRA ticket); not finding the same field as published in the literature (2.5 mG vs 15 mG); in the absence of Doppler tracking, looking at different runs of cvel to see how this affects things; ran a second test to confirm; looking at other tests to explore (e.g., W3OH).

Polarization: Taylor: Polarization calibration test.

  • Early update: test to look at transfer of primary to secondary for imaging, exploring calibration across the whole band (in this case 4.2 - 6.2 GHz; lose bits of the top due to known RFI). Recent setjy change to channelized flux density has helped. Still need a better polarization model for 3C286 (e.g., Q, U images) - needed for high fidelity broad band calibration. Another area that is lacking is in handling the curvature of the spectrum for calibrators - causes errors; AIPS has some handling of this (curvature over the band is handled). More later.

Spectral Line: Dhawan: Frequency offset tests (for Doppler tracking)

  • Wanted for Doppler tracking, want to see relative phase doesn't change as you sweep frequency; works for small frequency changes < 100 kHz, for most antennas (but not all!); also see effects of the antenna-based LO offset; both need to be solved; planning further tests in both lab and on sky (once we've crawled out from under the 3-bit work).

In preparation

  • Polarization: Multi-band leakage/PA (Myers, Marvil): Not a priority but still needed; keep on list.
  • Spectral Line: Bandpass stability,multi-band on 3C84 et al. (Goss/Chomiuk). 2/3rds through the data; report further in a week plus; data looks good.
  • Correlator: Report on 7 vs. 8 node observing (Krauss/Hunter)

Commissioning Test Planning

  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Delayed until Rick's return
  • Other: SysPower demonstration tests (in-out 1dB; long run (over a range of airmasses) on strong calibrator; Rick is developing the SBs. Follow-ups needed for Doppler tracking (phase jumps, sampling, etc).

RSRO/ECSO Reports

Project Band(s) Test Code Test Status Allocation Status (D) Notes JIRA
RSRO            
AA30 Q TRSR0003 - 0/51 Not started RSRO-1
AB1345 L TRSR0004 21May10 23/40 Partial; adequate for D RSRO-2
AB1346 K TRSR0005
TSPE0001
23Jun10 39/40 2 hour makeup pending for DnC RSRO-3
AC982 CKKaQ TRSR0002 28Jul10 30/30 Complete RSRO-4
AH1006 LCXKa TRSR0006 - 30/36 Partial RSRO-5
AK726 Ka TRSR0007 - 29.5/29.5 Complete RSRO-6
AL746 CKa TRSR0008 05Jul10 11/11 Complete RSRO-7
AM1014 C TRSR0009 22Jun10 0/0 C config RSRO-9
AS1015 C TRSR0010 21May10 4/4 Complete RSRO-11
AT374 C TRSR0011 25Jul10 0/0 C config RSRO-12
AV318 L TRSR0012 - 12/12 Complete; PI converted to OSRO RSRO-13
ECSO          
AW775 L TRSR0017 05Jul10 56/72 Partial -
AM1005 KKaQ TRSR0014 - 3/20 Partial (carry to C config) -
AM1038 Q TRSR0019 21May10 10/10 Complete -
AM1040 Ka TRSR0019 21May10 10/10 Complete -
AC992 L TRSR0018 26Jun10 4/4 Complete -
AO255 Ka TRSR0016 - 36/40 Complete -
AM1018 Ka TRSR0015 - 0/48 Needs test for DnC -
AB1353 Ka TRSR0013 - 0/51 Moved to C config -
  • Notes:
    • Current Residents
      • Laura Perez (POC: Claire Chandler); area(s): High frequency observing
      • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).
      • Volker Heesen (POC: Michael Rupen); area(s): polarimetry
      • Peter Hofner (POC: Juergen Ott); area(s): post-processing
      • Russ Taylor (POC: Rick Perley); area(s): polarimetry
      • NewSatoko Takahashi (POC: Juergen Ott); area(s): continuum, mosaicing; 10B-126 (AT380): A Millimeter Study of the Embedded Star-Cluster in the Orion Molecular Cloud.
        • NewHauyu Liu (POC: Juergen Ott); area(s): continuum, mosaicing
      • NewGreg Hallinan (POC: Kumar Golap); area(s):; 10B-209 (AH1020): Broadband Periodic Dynamic Spectra of Ultracool Dwarf Pulsars

Demo Science (Chandler/Condon): Project Codes: *TDEM*

Project Status Time Comments SB
Crab nebula Complete 10/10 Complete; L-band OSRO1/OSRO2  
IRC+10216 Ready (HF) 11/33 Partial; revisit in DnC/C  
M87 Ready (HF) 0/6 Ka Not done
M51 Ready (MR) 8/8 Complete  
Cygnus A Ready 3x6 1x10 possible; 12 OSRO hours obtained Withdrawn
M82 Complete 2/2 L band Complete
Herc A Ready 3/5 CX bands Partial

Commissioning Group updates

User Support Update (van Moorsel)

AOB

At the Cal-Im meeting in Dwingeloo 2 weeks ago, the ASTRON folks
advertised this image as a 'surprise' for NRAO attendees - and were
very pleased that Rick had given them these data.  Rick had told them
that the best that could be done with these data was a dynamic-range
of a few thousand to one.  Sarod-Yatawatta (ASTRON) ran it through a
multiscale-type imaging algorithm of his own, and produced a dynamic
range of 13000:1.  This was used to claim a superior algorithm
compared to those in any NRAO software.  Note however that it was not
known if anyone other than Rick/Bill at NRAO had imaged these data or
tried multi-scale/mem methods from either CASA or AIPS.

Misc Actions

  • [McMullin]: Splatalogue use for RFI line list.
  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin/Butler/Rupen/Owen]: Resolve discussion of requirements for DnC test of 4-band.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 08 Sep 2010

Topic attachments
I Attachment Action Size Date Who Comment
2GHZ-15X22-CAL.psps 2GHZ-15X22-CAL.ps manage 160 K 2010-09-14 - 16:35 UnknownUser  
2GHZ-15X22-RAW.psps 2GHZ-15X22-RAW.ps manage 159 K 2010-09-16 - 11:36 UnknownUser  
3B-LCP.psps 3B-LCP.ps manage 7 K 2010-09-16 - 11:35 UnknownUser  
8B-LCP.psps 8B-LCP.ps manage 7 K 2010-09-16 - 11:35 UnknownUser  
Topic revision: r6 - 2010-09-16, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback